# FAVAE-Effective Frequency Aware Latent Tokenizer

### v1

@@ Abstract
Latent generative models have shown remarkable progress in high-fidelity image synthesis, typically using a two-stage training process that involves compressing images into latent embeddings via learned tokenizers in the first stage. The quality of generation strongly depends on how expressive and well-optimized these latent embeddings are. While various methods have been proposed to learn effective latent representations, the reconstructed images often lack realism, particularly in textured regions with sharp transitions, due to loss of fine details governed by high frequencies. We conduct a detailed frequency decomposition of existing state-of-the-art (SOTA) latent tokenizers and show that conventional objectives inherently prioritize low-frequency reconstruction, often at the expense of high-frequency fidelity. Our analysis reveals these latent tokenizers exhibit a bias toward low-frequency information, when jointly optimized, leading to over-smoothed outputs and visual artifacts that diminish perceptual quality. To address this, we propose a wavelet-based, frequency-aware variational autoencoder (FA-VAE) framework that explicitly decouples the optimization of low- and high-frequency components. This decoupling enables improved reconstruction of fine textures while preserving global structure. Our approach bridges the fidelity gap in current latent tokenizers and emphasizes the importance of frequency-aware optimization for realistic image representation, with broader implications for applications in content creation, neural rendering, and medical imaging.
1 Introduction
Latent generative modeling Blattmann2023_VideoLDM ; Dong2024_GPLD3D ; Luo2024_LCM ; Zhou2024_LGD ; 10.1007/978-3-031-54605-1_10 ; medi20253dwaghierarchicalwaveletguidedautoregressive has emerged as a cornerstone of modern content creation, with recent advances demonstrating remarkable capabilities in synthesizing high-fidelity visual content. These models typically operate in compressed latent spaces learned via autoencoders such as VAEs kingma2013auto ; li2024autoregressive , where generation quality is directly influenced by the expressiveness of these latent embeddings. While increasing latent dimensionality can enhance representational power, this often leads to diminishing visual outputs and increased computational cost. VAVAE yao2025vavae addresses this by aligning the latent space with pre-trained vision foundation models to further improve convergence and generative realism. However, despite such improvements, generated outputs or even the reconstructed outputs often lack sharp textures and fine details, particularly in regions dominated by the high-frequency information such as text on images. This limits perceptual realism and results in overly smooth outputs. Prior works phung2023wavediff ; li2024fedsr ; guth2022wsgm have explored architectural and spectral enhancements to inject high-frequency signals, but a systematic frequency-level analysis of how these current latent tokenizers influence reconstruction quality, especially in differentiating low- vs. high-frequency information remains absent.
In this work, we conduct a comprehensive frequency-based analysis of the reconstruction behavior of state-of-the-art latent tokenizers used in generative pipelines yao2025vavae ; li2024autoregressive ; tian2024visual ; xie2024sana ; xie2025sana , focusing on latent diffusion and autoregressive models, which represent the current forefront of high-quality visual generation. These models typically follow a two-stage pipeline: (1) learning quantized or non-quantized latent embeddings using a latent tokenizer like VAE or its variant, and (2) training a probabilistic generative model in the latent embedding space. Our analysis reveals a consistent bias during the first stage of learning latent embeddings: while low-frequency components are on par well reconstructed, high-frequency signals such as textures are poorly preserved. This low frequency bias over high frequencies in optimization contributes significantly to perceptual degradation.
@@ 1 Introduction
Figure 1: Visual comparison of reconstructions. From left to right: example original image, VAVAE reconstruction, and our approach (FA-VAE). The red color highlighted regions emphasize areas rich in textures, edges, and text. Our method better preserves high-frequency details and sharp structures, resulting in reconstructions visually closer to the input.
@@
To address this fidelity gap, we propose a novel Frequency-aware VAE (FA-VAE) latent tokenizer that explicitly decouples and separately optimizes low- and high-frequency components using wavelet decomposition. Our method learns distinct latent embeddings for low-high frequency subbands and later fuses them into a unified latent space representation. This simple design allows for better preservation of both global structure and fine details. Figure 1 presents a visual comparison between our method and the recent state-of-the-art VAVAE, highlighting improved preservation of fine details and sharp structures.
@@
  • We provide a frequency-based analysis of latent tokenizers used in latent generative pipelines, showing that existing VAE variants disproportionately emphasize low-frequency components at the expense of fine detail.
@@
2 Method
2.1 Frequency Evaluation of Latent Embeddings
Latent tokenization models typically begin with a Variational Autoencoder (VAE) or its tokenizer variants, which learn compact embeddings $\mathbf{z}\in\mathcal{Z}$ from input data $\mathbf{x}\in\mathcal{X}$. Reconstruction quality is often measured by the pixel-level error.
$\mathcal{L}_{\text{rec}}=\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2},\quad\hat{\mathbf{x}}=D(E(\mathbf{x})),$
where $E$ and $D$ are the encoder and decoder.
@@ 2.1 Frequency Evaluation of Latent Embeddings
To study fidelity across frequency bands, we apply a discrete wavelet transform (DWT) with Haar filter to decompose signals into low- and high-frequency bands with one decomposition level:
@@
Then, we compute the frequency-aware reconstruction losses as follows:
@@
Beyond pixel errors, perceptual similarity is captured by LPIPS zhang2018unreasonable , and distributional alignment by reconstruction FID (rFID) heusel2017gans , computed from Inception features szegedy2016rethinking . Together, combining frequency-aware losses $(\mathcal{L}_{L},\mathcal{L}_{H})$, perceptual similarity (LPIPS), and rFID provides a more complete evaluation of reconstruction fidelity and expressiveness of the learned latent embeddings.
Figure 2: FA-VAE. The Overall pipeline of FA-VAE latent tokenizer.
2.2 Frequency-Aware VAE (FA-VAE)
The overview of FA-VAE is shown in Fig. 2. To enhance the fidelity of latent embeddings, we extend the standard VAE formulation by decoupling the input image into low- and high-frequency components, which are then learned independently. Given an input image $\mathbf{x}\in\mathcal{X}$, we apply a discrete wavelet transform $\mathcal{W}(\cdot)$ to obtain:
$\mathcal{W}(\mathbf{x})=(\mathbf{x}_{L},\mathbf{x}_{H}),$
where $\mathbf{x}_{L}$ and $\mathbf{x}_{H}$ denote the low- and high-frequency representations, respectively. We use the Haar filter for wavelet transformation, and apply the normalization strategy proposed in mulcahy1997image to normalize both low and high frequency components. We employ separate encoder–decoder pairs $(E_{L},D_{L})$ and $(E_{H},D_{H})$ to learn frequency-specific latent embeddings:
$\mathbf{z}_{L}=E_{L}(\mathbf{x}_{L}),\quad\mathbf{z}_{H}=E_{H}(\mathbf{x}_{H}).$
Low-Frequency Objective.
To learn low-frequency latent embeddings, we adopt a VA-VAE-style objective yao2025vavae , incorporating a vision foundation alignment loss $\mathcal{L}_{\text{VF}}^{L}$ yao2025vavae , adversarial regularization $\mathcal{L}_{\text{GAN}}^{L}$ inspired by VQGAN esser2021taming , and an additional perceptual loss $\mathcal{L}_{\text{LPIPS}}^{L}$ zhang2018unreasonable to improve visual quality. The total low-frequency objective is defined as:
where:
with $q(\mathbf{z}_{L}\mid\mathbf{x}_{L})$ being the encoder’s approximate posterior and $p(\mathbf{z}_{L})\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the standard gaussian prior. The vision foundation loss $\mathcal{L}_{\text{VF}}^{L}$ aligns the latent codes with the feature space of a pretrained foundation model (e.g., DINOv2 oquab2023dinov2 ). The adversarial loss $\mathcal{L}_{\text{GAN}}^{L}$ introduces a discriminator to distinguish real from reconstructed low-frequency inputs. Finally, the perceptual loss $\mathcal{L}_{\text{LPIPS}}^{L}$ encourages perceptual similarity between input and reconstruction using features from a pretrained Inception network szegedy2016rethinking . All these loss components are commonly used in modern VAE frameworks.
High-Frequency Objective.
At inference, both decoders reconstruct the respective frequency bands, and the final image is synthesized via inverse wavelet transform. This frequency-aware latent tokenization via FA-VAE yields more expressive embeddings by preserving details across the full spectrum of spatial frequencies.
3 Experiments
To evaluate the reconstruction performance of latent tokenizers in both the spatial and spectral (frequency) domains, we consider a range of widely adopted visual tokenizers commonly used in modern latent generative models. These tokenizers are based on different variants of autoencoding architectures, each aiming to learn compact and informative latent embedding representations of the input data distribution. For a comprehensive analysis, we include recent representative tokenizers based on standard Autoencoders chen2024deep , Variational Autoencoders (VAEs) rombach2022high ; li2024autoregressive ; lee2022autoregressive ; yu2024image ; yao2025vavae , and Vector Quantized Autoencoders (VQ-VAEs) yu2024image ; esser2021taming ; tian2024visual . We also include our proposed FA-VAE, which explicitly incorporates frequency-awareness into the latent embedding learning process. This selection enables a fair comparison across diverse tokenization strategies with varying latent dimensionalities (c) and latent feature resolutions (f) as shown in Table 1 with diverse tokenization configurations. We use the metrics discussed in method section 2.1 for our evaluation.
@@ 3 Experiments
Table 1: Quantitative comparison of latent tokenizers. Configurations: f = latent spatial resolution, c = latent dimensionality, v = vocabulary size (for quantized models). Lower is better.

### v2

# Missing Fine Details in Images: Last Seen in High Frequencies

@@ Abstract
Latent generative models have shown remarkable progress in high-fidelity image synthesis, typically using a two-stage training process that involves compressing images into latent embeddings via learned tokenizers in the first stage. The quality of generation strongly depends on how expressive and well-optimized these latent embeddings are. While various methods have been proposed to learn effective latent representations, generated images often lack realism, particularly in textured regions with sharp transitions, due to loss of fine details governed by high frequencies. We conduct a detailed frequency decomposition of existing state-of-the-art (SOTA) latent tokenizers and show that conventional objectives inherently prioritize low-frequency reconstruction, often at the expense of high-frequency fidelity. Our analysis reveals these latent tokenizers exhibit a bias toward low-frequency information during optimization, leading to over-smoothed outputs and visual artifacts that diminish perceptual quality. To address this, we propose a wavelet-based, frequency-aware variational autoencoder (FA-VAE) framework that explicitly decouples the optimization of low- and high-frequency components. This decoupling enables improved reconstruction of fine textures while preserving global structure. Moreover, we integrate our frequency-preserving latent embeddings into a SOTA latent diffusion model, resulting in sharper and more realistic image generation. Our approach bridges the fidelity gap in current latent tokenizers and emphasizes the importance of frequency-aware optimization for realistic image synthesis, with broader implications for applications in content creation, neural rendering, and medical imaging.
Introduction
@@
Figure 1: Visual comparison of reconstructions. From left to right: original image, VAVAE reconstruction, and our approach. The highlighted regions emphasize areas rich in textures, edges, and text. Our method better preserves high-frequency details and sharp structures, resulting in reconstructions visually closer to the input.
Latent generative modeling (Blattmann et al. 2023; Dong et al. 2024; Luo et al. 2024; Zhou, Wang, and Zhang 2024) has emerged as a cornerstone of modern content creation, with recent advances demonstrating remarkable capabilities in synthesizing high-fidelity visual content. These models typically operate in compressed latent spaces learned via autoencoders such as VAEs (Kingma and Welling 2013; Li et al. 2024a), where generation quality is directly influenced by the expressiveness of the latent embeddings. While increasing latent dimensionality can enhance representational power, this often leads to diminishing visual outputs and increased computational cost. VAVAE (Yao, Yang, and Wang 2025) addresses this by aligning the latent space with pre-trained vision foundation models to improve convergence and generative realism. However, despite such improvements, generated outputs often lack sharp textures and fine details, particularly in regions dominated by the high-frequency information such as text on images. This limits perceptual realism and results in overly smooth reconstructions. Prior works (Phung, Dao, and Tran 2023; Li et al. 2024b; Guth et al. 2022) have explored architectural and spectral enhancements to inject high-frequency signals, but a systematic frequency-level analysis of how these current latent embeddings influence reconstruction quality, especially in differentiating low- vs. high-frequency information remains absent.
Figure 2: Residual power spectra averaged over 50k ImageNet validation images, comparing reconstruction errors (input minus reconstruction) of VAVAE and our method. The log-scaled spectra show that VAVAE exhibits higher residual energy across the frequency spectrum, particularly in high-frequency regions. In contrast, our method significantly reduces reconstruction residual energy, indicating better preservation of fine details and textures.
In this work, we conduct a comprehensive frequency-based analysis of the reconstruction behavior of state-of-the-art latent embeddings used in generative pipelines (Yao, Yang, and Wang 2025; Li et al. 2024a; Tian et al. 2024; Xie et al. 2024, 2025), focusing on latent diffusion and autoregressive models, which represent the current forefront of high-quality visual generation. These models typically follow a two-stage pipeline: (1) learning quantized or non-quantized latent embeddings using a VAE variant, and (2) training a probabilistic generative model in the latent space. Our analysis reveals a consistent bias: while low-frequency components are well reconstructed, high-frequency signals such as textures and fine structures are poorly preserved. This low frequency bias over high frequencies in optimization contributes significantly to perceptual degradation.
@@
To address this fidelity gap, we propose a novel Frequency-aware VAE (FA-VAE) framework that explicitly decouples and separately optimizes low- and high-frequency components via wavelet decomposition. Our method learns distinct latent embeddings for frequency subbands and later fuses them into a unified latent space representation. This simple design allows for better preservation of both global structure and fine details. Figure 1 presents a visual comparison between our method and the recent state-of-the-art VAVAE, highlighting improved preservation of fine details and sharp structures. Complementarily, Figure 2 quantifies reconstruction quality in the frequency domain using the average residual power spectrum of reconstruction errors computed with respect to input over 50K ImageNet (Deng et al. 2009) validation images. Our method consistently exhibits lower residual energy across both low (middle of spectrum) and high (corners of the spectrum) frequency bands, with a particularly notable reduction in high-frequency at corners of the power spectrum. This indicates more faithful reconstructions over the frequency spectrum, directly contributing to the enhanced perceptual quality observed in Figure 1. Finally, we integrate our frequency aware latent embeddings into a lightweight latent diffusion model (LightningDiT, based on VAVAE), to showcase our frequency-aware embeddings enable sharper, more realistic generations.
@@
  • We provide a frequency-based analysis of latent embeddings used in latent generative pipelines, showing that existing VAE variants disproportionately emphasize low-frequency components at the expense of fine detail.
@@
  • We integrate FA-VAE embedding space into a diffusion pipeline (LightningDiT), demonstrating superior generation quality and perceptual fidelity across ImageNet, underscoring the role of frequency-aware modeling in high-fidelity synthesis.
Related Work
Figure 3: Overview of our proposed frequency-aware VAE framework (FA-VAE). An input image is decomposed into low- and high-frequency representations using a wavelet transform. Each frequency band is decoupled, encoded and decoded separately to learn dedicated latent embeddings, which are then coupled and passed through an inverse wavelet transform to reconstruct the image. These enriched embeddings are subsequently used in a latent diffusion model to improve generation fidelity, particularly preserving fine details.
Latent Embeddings for Image Generation
Latent embeddings or tokenizers are widely used to represent compressed, high-dimensional visual data and enable efficient generation. Encoder–Decoder frameworks such as VAEs (Kingma and Welling 2013) and their variants form the foundation for learning these latent representations, which are vital to many latent generative models. Notably, Latent Diffusion Models (LDMs) (Rombach et al. 2022) operate over learned latent embeddings to facilitate high-quality synthesis, highlighting the importance of effective latent space modeling. Discrete tokenization techniques, such as VQ-VAE (Van Den Oord, Vinyals et al. 2017; Medi et al. 2024b), convert images h code sequences that can be effectively modeled by transformers. VQGAN (Esser et al. 2021) further improves this framework by incorporating adversarial losses to learn perceptually meaningful tokens, while RQ-VAE (Lee et al. 2022) enhances expressiveness of embeddings via stacked residual quantizers. Recent methods like TiTok (Yu et al. 2024a), FlexTok (Bachmann et al. 2025), and OneDPiece (Miwa et al. 2025) explore compact and adaptive tokenization strategies, aiming for better compression and controllable generation quality.
To enhance reconstruction quality of latent embeddings, hierarchical and regularized latent models have been proposed. NVAE (Vahdat and Kautz 2020) introduces spectral regularization within residual hierarchies for stable training, while HQ-VAE (Takida et al. 2023) and VAR (Tian et al. 2024) use multi-scale vector quantization to better model local and global structure. EQ-VAE (Kouzelis et al. 2025) imposes equivariance constraints for robustness to geometric changes. Additionally, frequency-aware losses such as fourier-based regularization (Björk, Myhre, and Johansen 2022) and scale-equivariant decoding (Skorokhodov et al. 2025) have been used to control the spectral properties of the latent reconstructions. A parallel direction focuses on aligning latent spaces with external vision encoders. REPA (Yu et al. 2024b) introduces denoising-time regularization based on foundation model features to accelerate convergence. VA-VAE (Yao, Yang, and Wang 2025) builds on this by jointly aligning tokenizer latents with pretrained features and optimizing for downstream generation quality. ReaLS (Xu et al. 2025) introduces semantic priors directly into the latent space to improve representation capacity. Our work builds on VA-VAE but identifies that such alignment alone does not sufficiently preserve high frequency details. We propose an explicit decoupling and optimization of low- and high-frequencies during tokenization, leading to improved fidelity.
High-Frequency Information in Image Synthesis
High-frequency content is essential for perceptual realism but challenging to reconstruct, particularly for diffusion-based models. Prior analyses show that frequency components are not reconstructed uniformly across the denoising trajectory (Yang et al. 2023), resulting in a loss of textures and sharp structures. To address this, various frequency-aware approaches have been developed across generative paradigms. In GANs, SWAGAN (Gal et al. 2021) and WaveFill (Yu et al. 2021) integrate wavelet transforms into their architectures to improve generation fidelity. Focal frequency loss (Jiang et al. 2021) and WINE (Kim, Moon, and Park 2025) address frequency imbalance by reweighting reconstruction errors and correcting low-frequency bias in inversion, respectively.
Diffusion models have incorporated frequency priors at both training and inference time. WaveDiff (Phung, Dao, and Tran 2023), Spatial-Frequency UNet (Yuan et al. 2023), and Spectral Diffusion (Yang et al. 2023) integrate wavelet-domain frequency components or gating mechanisms into the network to better optimize the spectral content. Training-free techniques like DiffuseHigh (Kim et al. 2025) and frequency aggregation (Qian et al. 2024) inject wavelet-based frequency priors during inference for improved detail preservation. Additional strategies include replacing convolutions with filtered variants in FouriScale (Huang et al. 2024), and autoregressive frequency modeling in NFIG (Huang et al. 2025). Adaptations also exist for domain-specific tasks such as high-resolution (Zhang et al. 2025), underwater (Zhao et al. 2024), or low-light image generation (Jiang et al. 2023). Beyond images, wavelet representations have further proven useful in 3D generation with diffusion and autoregressive models (Hui et al. 2024; Medi et al. 2024a), suggesting the utility of wavelet decompositions as a general tool for capturing fine-scale structure across visual modalities. While effective, these approaches typically introduce additional architectural components or post-processing steps to compensate for frequency degradation. In contrast, our method addresses the root cause by improving frequency preservation at the latent representation level. By modeling and optimizing frequency subbands (low and high) separately during tokenization, we improve the frequency aware expressivity in the latent embeddings. These latent embeddings improve both reconstruction fidelity and downstream synthesis quality when integrated with latent generative pipeline without requiring any architectural changes at generation time.
Method
In this section, we present three key contributions. First, we conduct a detailed frequency analysis of latent embeddings commonly used in high-fidelity image generation. Second, we propose FA-VAE, a novel framework that decouples and independently optimizes low- and high-frequency components of the embeddings. Third, we introduce a simple fusion strategy to unify both components into a single latent representation, which is integrated into the LightningDiT pipeline (Yao, Yang, and Wang 2025), resulting in improved image generation via frequency-aware latent modeling. An overview of our framework is shown in Figure 3.
Frequency Evaluation of Latent Embeddings
In most latent generative models, the first stage of training includes Variational Autoencoders (VAEs) or their other visual tokenizer variants, aiming to learn compact latent embeddings $\mathbf{z}\in\mathcal{Z}$ that capture the underlying structure of the input data $\mathbf{x}\in\mathcal{X}$. A common measure of representational quality is the reconstruction error computed via the mean squared error (MSE) or $\ell_{2}$-loss:
$\mathcal{L}_{\text{rec}}=\|\mathbf{x}-\hat{\mathbf{x}}\|_{2}^{2},$
where $\hat{\mathbf{x}}=D(E(\mathbf{x}))$ is the reconstruction obtained by encoding $\mathbf{x}$ to a latent representation $\mathbf{z}=E(\mathbf{x})$, followed by decoding with decoder $D$ to obtain reconstruction $\hat{\mathbf{x}}$.
@@
Similarly, To assess the fidelity of reconstructions across low and high frequency bands, we employ a discrete wavelet transform (DWT) using Haar filter to decompose both the input and its reconstruction into respective low- and high-frequency representations. Let $\mathcal{W}(\cdot)$ denote the wavelet decomposition operator:
@@
where $\mathbf{x}_{L}$ and $\hat{\mathbf{x}}_{L}$ correspond to the low-frequency (approximation) representations, and $\mathbf{x}_{H}$, $\hat{\mathbf{x}}_{H}$ correspond to the high-frequency (detail) representations. We then compute frequency-specific reconstruction losses:
@@
This evaluation provides insight into the frequency fidelity of latent embeddings, measuring their ability to capture both low and high frequency information from input representation.
[…]
$\text{rFID}=\|\mu_{\mathbf{x}}-\mu_{\hat{\mathbf{x}}}\|_{2}^{2}+\text{Tr}\left(\Sigma_{\mathbf{x}}+\Sigma_{\hat{\mathbf{x}}}-2\left(\Sigma_{\mathbf{x}}\Sigma_{\hat{\mathbf{x}}}\right)^{1/2}\right),$
[…]
Together, the frequency-aware losses $(\mathcal{L}_{L},\mathcal{L}_{H})$, perceptual similarity $\mathcal{L}_{\text{LPIPS}}$, and distributional alignment via rFID provide a comprehensive evaluation of the reconstruction quality and frequency fidelity of latent embeddings.
@@
Table 1: Quantitative comparison of various latent tokenizers based on their reconstruction quality across multiple metrics. Tokenizer configurations are denoted as: f : latent spatial resolution, c: latent dimensionality, and v: vocabulary size in case of quantized models. Lower values indicate better performance.
Frequency-Aware VAE (FA-VAE)
To enhance the fidelity of latent embeddings, we extend the standard VAE formulation by decoupling the input image into low- and high-frequency components, which are then learned independently. Given an input image $\mathbf{x}\in\mathcal{X}$, we apply a discrete wavelet transform $\mathcal{W}(\cdot)$ to obtain:
$\mathcal{W}(\mathbf{x})=(\mathbf{x}_{L},\mathbf{x}_{H}),$
where $\mathbf{x}_{L}$ and $\mathbf{x}_{H}$ denote the low- and high-frequency representations, respectively. We use the Haar filter for wavelet transformation, and apply the normalization strategy proposed in (Mulcahy 1997) to normalize both low and high frequency components. We employ separate encoder–decoder pairs $(E_{L},D_{L})$ and $(E_{H},D_{H})$ to learn frequency-specific latent embeddings:
$\mathbf{z}_{L}=E_{L}(\mathbf{x}_{L}),\quad\mathbf{z}_{H}=E_{H}(\mathbf{x}_{H}).$
Low-Frequency Objective.
To learn low-frequency latent embeddings, we adopt a VA-VAE-style objective (Yao, Yang, and Wang 2025), incorporating a vision foundation alignment loss $\mathcal{L}_{\text{VF}}^{L}$ (Yao, Yang, and Wang 2025), adversarial regularization $\mathcal{L}_{\text{GAN}}^{L}$ inspired by VQGAN (Esser et al. 2021), and an additional perceptual loss $\mathcal{L}_{\text{LPIPS}}^{L}$ (Zhang et al. 2018) to improve visual quality. The total low-frequency objective is defined as:
where:
with $q(\mathbf{z}_{L}\mid\mathbf{x}_{L})$ being the encoder’s approximate posterior and $p(\mathbf{z}_{L})\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, the standard Gaussian prior.
The vision foundation loss $\mathcal{L}_{\text{VF}}^{L}$ aligns the latent codes with the feature space of a pretrained foundation model (e.g., DINOv2 (Oquab et al. 2023)). The adversarial loss $\mathcal{L}_{\text{GAN}}^{L}$ introduces a discriminator to distinguish real from reconstructed low-frequency inputs. Finally, the perceptual loss $\mathcal{L}_{\text{LPIPS}}^{L}$ encourages perceptual similarity between input and reconstruction using features from a pretrained Inception network (Szegedy et al. 2016). All these loss components are commonly used in modern VAE frameworks.
High-Frequency Objective.
Inference.
At inference time, both decoders are used to reconstruct the respective frequency bands, and the final image is synthesized via inverse wavelet transform:
This frequency-aware latent representation learning via FA-VAE yields more expressive embeddings by preserving both coarse structure and fine details across the full spectrum of spatial frequencies.
Latent Fusion for Generative Modeling
To enable frequency-aware generation, we integrate the proposed FA-VAE into the recent latent diffusion model LightningDiT (Yao, Yang, and Wang 2025) by considering the learned low-frequency embedding $\mathbf{z}_{L}\sim q(\mathbf{z}\mid\mathbf{x})$ as the primary latent representation. To enrich this representation, we fuse it with a complementary high-frequency latent code $\mathbf{z}_{H}$, which is aligned to enhance the representational capacity of $\mathbf{z}_{L}$. The two embeddings are combined using a lightweight fusion function:
$\tilde{\mathbf{z}}=\mathcal{F}(\mathbf{z}_{L},\mathbf{z}_{H}),$
where $\mathcal{F}$ denotes a simple concatenation operation that preserves both spatial and semantic alignment between the frequency bands. The resulting fused latent embedding $\tilde{\mathbf{z}}$ retains global structural coherence from $\mathbf{z}_{L}$, while incorporating fine-scale details through $\mathbf{z}_{H}$. This design is fully modular and plug-and-play, allowing seamless integration into existing generative architectures without requiring modifications to the underlying model structure.
Experiments
In this section, we present comprehensive evaluations to support our proposed contributions. We first assess the reconstruction performance of various latent embedding tokenizers, including our FA-VAE, in both spatial and spectral domains. To evaluate their generative utility, we integrate the embeddings into standard image synthesis pipelines, comparing FA-VAE combined with LightningDiT (Yao, Yang, and Wang 2025). All experiments are conducted on ImageNet-1K (Deng et al. 2009) and use 50k Imagenet Validation samples for quantitative evaluations. Beyond visual quality, we also examine embedding fairness across challenging ImageNet categories, comparing against VAVAE, the best and recent baseline. Results show that our frequency-aware embeddings improve reconstruction fidelity and promote a more fair latent space, especially for structurally complex classes. Further implementation details of our experiments are provided in Appendix A.2.
Performance Evaluation of Latent Embeddings
To evaluate the reconstruction performance of latent embeddings in both the spatial and spectral (frequency) domains, we consider a range of widely adopted visual tokenizers commonly used in modern latent generative models. These tokenizers are based on different variants of autoencoding architectures, each aiming to learn compact and informative latent embedding representations of the input data distribution. For a comprehensive analysis, we include recent representative tokenizers based on standard Autoencoders (Chen et al. 2024), Variational Autoencoders (VAEs) (Rombach et al. 2022; Li et al. 2024a; Lee et al. 2022; Yu et al. 2024a; Yao, Yang, and Wang 2025), and Vector Quantized Autoencoders (VQ-VAEs) (Yu et al. 2024a; Esser et al. 2021; Tian et al. 2024). We also include our proposed FA-VAE, which explicitly incorporates frequency-awareness into the latent embedding learning process. This selection enables a fair comparison across diverse tokenization strategies with varying latent dimensionalities (c) and latent feature resolutions (f) as shown in Table 1 with diverse tokenization configurations. We use the metrics discussed in method section for our evaluation.
Figure 4: Generated Visualization of our proposed FA-VAE together with LightningDiT-XL trained on ImageNet 256 × 256 resolution.
Method Tokenizer Epochs #Params Generation w/o CFG Generation w/ CFG
gFID sFID IS Pre. Rec. gFID sFID IS Pre. Rec.
Autoregressive (AR)
MaskGIT VQ-VAE 555 227M 6.18 - 182.1 0.80 0.51 - - - - -
LlamaGen VQ-VAE 300 3.1B 9.38 8.24 112.9 0.69 0.67 2.18 5.97 263.3 0.81 0.58
VAR MS-VQ-VAE 350 2.0B - - - - - 1.80 - 365.4 0.83 0.57
MagViT-v2 VQ-VAE 1080 307M 3.65 - 200.5 - - 1.78 - 319.4 - -
MAR KL-VAE 800 945M 2.35 - 227.8 0.79 0.62 1.55 - 303.7 0.81 0.62
Latent Diffusion Models
MaskDiT SD-VAE 1600 675M 5.69 10.34 177.9 0.74 0.60 2.28 5.67 276.6 0.80 0.61
DiT SD-VAE 1400 675M 9.62 6.85 121.5 0.67 0.67 2.27 4.60 278.2 0.83 0.57
SiT SD-VAE 1400 675M 8.61 6.32 131.7 0.68 0.67 2.06 4.50 270.3 0.82 0.59
FasterDiT SD-VAE 400 675M 7.91 5.45 131.3 0.67 0.69 2.03 4.63 264.0 0.81 0.60
MDT SD-VAE 1300 675M 6.23 5.23 143.0 0.71 0.65 1.79 4.57 283.0 0.81 0.61
MDTv2 SD-VAE 1080 675M - - - - - 1.58 4.52 314.7 0.79 0.65
REPA SD-VAE 800 675M 5.90 - - - - 1.42 4.70 305.7 0.80 0.65
LightningDiT VA-VAE 64 675M 5.14 4.22 130.2 0.76 0.62 2.11 4.16 252.3 0.81 0.58
LightningDiT FA-VAE 64 675M 3.24 4.09 193.7 0.83 0.69 1.32 4.07 317.4 0.83 0.65
Table 2: Comparision of generation performance of autoregressive and latent diffusion models with and without classifier-free guidance (CFG) vs LightningDiT with our FA-VAE tokenizer. Missing values are indicated by -. We report metrics only available in the respective works.
Model Rec. LF HF LPIPS rFID
KL-VAE 0.1146 0.0645 0.1313 0.2151 0.7820
MS-VQ-VAE 0.0238 0.0705 0.0082 0.2331 0.8427
VA-VAE 0.0125 0.0263 0.0098 0.1071 0.6509
FA-VAE 0.0044 0.0114 0.0020 0.0940 0.4156
Table 3: Reconstruction comparison of tokenizers with input wavelet representation instead of input pixel representation. Rec.: Reconstruction Loss, LF: Low Frequency Loss, HF: High Frequency Loss. Lower is better.